# Z-Dip: a validated generalization of the Dip Test

### v1

@@ 6 Conclusions
This work introduced the Z-Dip, a standardized reformulation of Hartigan’s Dip Test that enables consistent and interpretable multimodality assessment across sample sizes. By expressing the Dip in standardized units relative to its null distribution, the Z-Dip transforms a size-dependent hypothesis test into a scale-free descriptive measure. The proposed framework retains the theoretical foundation and computational efficiency of the original Dip Test, while eliminating the need for lookup tables and enabling direct comparison of multimodality strength across datasets.
Validation on synthetic samples and empirical data demonstrate that the Z-Dip preserves the inferential behavior of the Dip Test, achieving virtually perfect correspondence in significance decisions. At the same time, the standardized scale provides additional interpretability, allowing researchers to quantify how strongly a distribution deviates from unimodality.
Finally, we proposed a downsampling-based correction to stabilize Z-Dip values in very large samples, enabling robust comparisons across datasets of varying size. Together, these contributions establish the Z-Dip as a practical, transparent, and computationally efficient tool for detecting and quantifying multimodality. We believe that the availability of reference tables and open-source implementations will facilitate its adoption in applications in a wide range of fields.
Acknowledgements
We thank Parnasso Nesso for their precious inspiration and contribution.
Supplementary Information
@@ Acknowledgements
Figure S1: Scaling of Z-Dip as NN grows. The average Z-Dip values were computed over 100 samples coming from the same underlying distribution, for each NN.

### v2

@@ 6 Conclusions
This work introduced the Z-Dip, a standardized reformulation of Hartigans’ Dip Test that enables consistent and interpretable modality assessment across sample sizes. By expressing the Dip in standardized units relative to its null distribution, the Z-Dip transforms a size-dependent hypothesis test into a scale-free descriptive measure. The proposed framework retains the theoretical foundation and computational efficiency of the original Dip Test, while eliminating the need for lookup tables and enabling direct comparison of multimodality strength across datasets.
Validation on synthetic samples and empirical data demonstrate that the Z-Dip preserves the inferential behavior of the Dip Test, achieving virtually perfect correspondence in significance decisions. At the same time, the standardized scale provides additional interpretability, allowing researchers to quantify how strongly a distribution deviates from unimodality. Finally, we proposed a downsampling-based correction to stabilize Z-Dip values in very large samples, enabling robust comparisons across datasets of varying size.
Together, these contributions establish the Z-Dip as a practical, transparent, and computationally efficient tool for detecting and quantifying multimodality. We believe that the availability of reference tables and open-source implementations will facilitate its adoption in applications in a wide range of fields.
Data and code availability
An open-source R implementation of the Z-Dip is available at https://github.com/EdoardoDima/zdip. The package can be installed locally by following the instructions provided in the repository. The data used in the empirical validation are available at https://osf.io/vh9px/. Further details on the inference of users’ political leaning in the YouTube videos analyzed can be found in [21].
CRediT authorship contribution statement
Edoardo Di Martino: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Visualization, Writing - Original Draft, Writing - Review & Editing, Matteo Cinelli: Conceptualization, Supervision, Visualization, Project administration, Writing - Review & Editing, Roy Cerqueti: Conceptualization, Supervision, Project administration, Writing - Review & Editing,
Supplementary Information for Z-Dip: a validated generalization of the Dip Test for data modality assessment
Rescaling the Z-Dip to a fixed scale
As mentioned in the main manuscript, the Z-Dip expresses how strongly the observed distribution deviates from unimodality in terms of standard deviation relative to the null expectation. While intuitive, this measure remains unbounded by definition, and we acknowledge how limiting the scores to a certain interval could make the measure more interpretable in certain applications. To do this, a number of squashing functions could be used to map the scores in a certain interval: we show as an example a sigmoid-based function, which remaps scores in the interval [1,+1][-1,+1]:
$\sigma(z)=\frac{2}{1+e^{-\alpha z}}-1,$
for which we propose as a standard $\alpha=0.595$, the choice of which neatly remaps the threshold $z=1.85$ to an intuitive $\sigma(z)=0.5$. Figure S1 shows the distribution of the rescaled scores computed on 530,000530{,}000 uniform samples with sample size ranging from $N=4$ to N=72, [^0] N=72{,}000.
Figure S1: Distributions of Z-Dip scores after applying a sigmoid transformation with $\alpha=0.595$. The dashed line shows the rescaled significance threshold $\sigma(z)=0.5$.
@@
Figure S2: Scaling of Z-Dip as $N$ grows, for explicitly multimodal samples. The average Z-Dip values were computed over 100 samples coming from the same underlying distribution, for each NN.